# Can Few-shot Work in Long-Context? Recycling the Context to Generate Demonstrations

### v3

@@ 4 Experiments
Gemini Nano (vanilla) 41.6 10.8 72.2 66.8 55.4 21.3 59.6 5.2
   Zero-shot + Evidence Retrieval 56.5 32.0 82.4 82.4 56.4 56.7 60.5 25.2
   DoubleDipper 62.1 40.6 86.6 95.4 56.2 65.1 60.5 30.4
Gemma 2 2B (v2) (vanilla) 38.6 8.9 71.8 68.6 51.2 13.3 49.6 6.5
   Zero-shot + Evidence Retrieval 42.0 22.3 66.8 70.6 40.2 30.6 47.6 16.2
   DoubleDipper 49.5 23.7 85.8 81.6 50.0 39.9 46.7 18.8
@@
Mixtral 7x8B (v0.1) (vanilla) 42.6 13.7 73.0 66.2 51.0 18.2 67.7 8.4
   Zero-shot + Evidence Retrieval 47.4 18.8 81.8 73.6 50.6 26.3 67.9 13.1
   DoubleDipper 52.2 22.3 91.8 86.0 47.8 35.1 66.6 16.0
@@
Models
We apply DoubleDipper to a variety of models, both commercial and open-source. The commercial models include Gemini 1.5 Pro, Gemini 1.5 Flash and Gemini 1.0 Nano (Reid et al., 2024). The open-source models we tested are Llama 3.1 8B (Dubey et al., 2024), Gemma 2B (v2) and Gemma 9B (v2) (Riviere et al., 2024), Mistral-7B-Instruct (v0.2) (Jiang et al., 2023) and Mixtral-8x7B-Instruct (v0.1) (Jiang et al., 2024). Details about models’ size and context window are shown in Table 1.
Baselines
We compare DoubleDipper to the vanilla baseline, which takes as input the entire context $D$ and the query $q$ and generates only the answer $a$. We also compare to Zero-shot + Evidence Retrieval, which prompts the model in a zero-shot setting to identify the relevant passage(s) before generating the answer.
Evaluation
We evaluate each dataset with the original evaluation metrics. Namely, we report Accuracy for Lost-in-the-middle (Liu et al., 2023) and FLenQA (Levy et al., 2024), and Token F1 for HotPotQA (Yang et al., 2018), 2Wiki (Ho et al., 2020) and MuSique (Trivedi et al., 2021).
Implementation Details
We randomly select three passages from the input (see Section 6 for an analysis of the number of self-generated demonstrations on the performance), each containing at least two sentences, and ask the model to generate a single QA pair for each passage (see Appendix A for the exact prompt). For all experiments, including few-shot generation and question-answering, we use a temperature setting of 0.
5 Results
@@
Table 3 presents the QA performance of the baseline, Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set. The results first show that prompting models to explicitly identify the relevant paragraphs before generating the answer (Zero-shot + Evidence Retrieval) leads to a performance improvement of 9.7 points on average across models over the vanilla baseline. DoubleDipper offers an additional substantial boost of 6.3 points for all models on average, culminating in a overall improvement of 16 absolute points over the vanilla baseline. Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it always surpasses the baseline in multi-hop QA datasets (HotPotQA, 2Wiki and MuSique). Likewise, DoubleDipper outperforms the baseline also on the FLenQA datasets (PIR, MonoRel and SRT), which involve synthetic True/False questions although the demonstrations in DoubleDipper are typically simple factoid questions.
@@
Gemini Nano Zero-shot + Evidence Retrieval 68.1 76.3 77.6 71.6 66.0 74.1 55.8 55.2
DoubleDipper 69.9 76.6 86.6 74.5 59.8 76.0 59.2 56.7
Gemma 2B Zero-shot + Evidence Retrieval 39.1 57.6 49.4 44.3 18.9 53.1 14.0 36.4
DoubleDipper 45.6 56.5 60.7 57.5 13.1 58.1 33.1 39.9
Gemma 9B Zero-shot + Evidence Retrieval 61.9 76.7 69.3 60.3 43.2 76.5 51.5 55.9
@@
Figure 4: Performance (accuracy) of Gemma 2 9B, Mixtral 8x7B and Gemini Flash with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
Result 2: Learning to retrieve the evidence(s) with DoubleDipper is more effective in commercial models than in open source.
Table 4 presents the performance of the supporting paragraphs prediction for the Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set. For all commercial models and Gemma 2B, DoubleDipper predicts better the supporting paragraphs than in the zero-shot setting (+2.6 F1 for Gemini Pro, +3.4 F1 for Gemini Flash, +1.8 F1 for Gemini Nano and +6.5 F1 for Gemma 2B). On the other hand, DoubleDipper slightly hurts the performance of common open source models (e.g., -2.8 F1 for Mistral and -0.3 F1 for Llama 3.1). This difference is because DoubleDipper’s demonstrations are based on a single paragraph, where commercial models can generalize better and predict multiple evidences. Indeed, Gemini Pro predicts on average 2 evidences for the datasets 2Wiki, MuSique, PIR, MonoRel, SRT and HotPotQA whereas Gemma 9B predicts 1.2 evidences for each instance. In fact, the only dataset with a single evidence is “Lost” and almost all models highly benefit from DoubleDipper (e.g, +11.9 F1 for Llama 3.1, +12.2 F1 for Mistral 7B and +9.8 for Mixtral 7x8B).
@@
Following Liu et al. (2023), Figure 4 shows the performance of Gemma 2 9B, Mixtral 8x7B and Gemini Flash for both the baseline and DoubleDipper on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer. See Appendix B for the performance curve of the other tested models, which show similar trends.
@@ 6 Ablation Studies
Gemma 2B Self Gemini Pro
Gemini Pro 70.4 71.6 71.6
Gemini Nano 62.1 61.7 62.8
Gemini Flash 66.1 67.5 68.1
Gemma 2B 49.5 49.5 51.0
Gemma 9B 61.2 62.0 63.5
Llama 3.1 59.9 60.1 61.5
Mistral v0.3 51.0 49.8 52.6
Mixtral 52.2 49.8 54.4
Table 5: Average performance of DoubleDipper with different models for generating the demonstrations. See Appendix C.2 for the results on each evaluation dataset.
How many examples are needed?
In Table 6, we explore the impact of varying $k$, the number of self-generated few-shot examples in DoubleDipper to 1, 3, 5, and 10. On average, a single demonstration already provides an improvement over the baseline. Adding 3 demonstrations adds another boost of 2 points, while increasing the number of demonstrations to 5 and 10 leads to a marginal improvement. This finding is in line with previous work (Brown et al., 2020a; Min et al., 2022b). We conclude that a small number of examples carries most of the benefit with our method, but given additional computation budget, adding more examples does carry additional minor benefit.
@@
Table 6: Average performance on our evaluation set with various numbers of self-generated few shot demonstrations ($k$) in DoubleDipper. See Appendix C.1 for the results on each evaluation dataset.
@@
To understand the impact of the default chosen model (Gemma 2 2B) for generating the demonstrations, we conducted two additional experiments. The first experiment is Self in which we use the same model for generating the demonstrations and for answering the original question. In the second experiment, we generate the demonstrations with the best LLM used in our experiments, namely Gemini Pro. The average results are reported in Table 5 and the performance for each evaluation dataset is presented in Appendix C.2. The results show that generating the demonstrations with Gemma 2 or Self achieves similar performance, while Gemini Pro leads to a consistent increase in performance across models, indicating that future better models can improve further the performance.
As mentioned in Section 3, DoubleDipper promotes also efficiency by adding to the original prompt only a few extra tokens, leading to a significantly cheaper inference than the traditional ICL.
DoubleDipper without identification of supporting paragraphs
To ablate the second principle in DoubleDipper, which is the explicit identification of the supporting paragraphs before generating the answer, we prompt the open source models with self-generated few shot examples that comprise only question-answer pairs (without instructing the model to retrieve the relevant passage(s)). For all models, the average QA performance of DoubleDipper without the evidence is lower than DoubleDipper with evidence identification. When averaging results across models and datasets, we report a substantial F1 drop from 54.8 to 46.6. Full results are available in Appendix C.3 in Table 9.
@@ 7 Conclusion
We develop DoubleDipper, a straightforward method for enhancing the performance of Question Answering with long context and providing attribution to the relevant paragraph(s) in the input. By recycling the input context to generate the few shot examples, each demonstration includes solely a question, an answer and a pointer to the relevant paragraph, without a separate context, thus effectively addressing the challenging of In-Context-Learning with long context. Experimental results show that our approach substantially ourperforms the baselines in various QA settings, including distractor passages in the input, True/False questions and multi-hop QA.
@@ 8 Limitations
One notable limitation of our approach is the extended inference time required for generating question-answer pairs. Future research could mitigate this issue by developing smaller, specialized models specifically tailored for QA generation.
Additionally, our evaluation set is constrained to instances that are solely in English and range between 1,000 to 4,000 tokens. Expanding the diversity of languages and token ranges could enhance the robustness and applicability of our findings.
Lastly, although we employ a strategy of randomly sampling $k$ paragraphs from the input to ensure the model engages with varied segments of the text, we did not optimize the selection of these paragraphs. Future work could explore more strategic methods for paragraph selection to potentially enhance the efficacy and relevance of the generated examples.
Appendix A Prompts
@@
Appendix B Lost-in-the-middle
@@
Figure 6: Performance (accuracy) of Gemma 2 2B, Mistral 7B, Gemini Nano and Gemini Pro with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023) according to the position of the document that contains the answer.
Appendix C Analysis
C.1 Impact of the Number of Demonstrations in DoubleDipper
@@
C.2 Impact of the few-shot generator
@@
C.3 Impact of the identification of supporting paragraphs in the QA generation

### v4

@@ 5 Results
Gemma 2 27B (v2) (vanilla) 48.0 11.2 85.2 79.2 63.8 14.3 62.4 19.9
   Zero-shot + Evidence Retrieval 59.7 34.3 90.6 87.4 59.0 51.7 65.2 29.8
   DoubleDipper 64.2 42.0 92.0 96.4 58.6 64.2 62.5 33.9
@@
Llama 3.1 70B (vanilla) 67.5 46.5 93.2 95.6 83.2 57.0 69.6 27.6
   Zero-shot + Evidence Retrieval 71.0 57.6 97.8 97.6 81.0 62.0 71.5 29.5
   DoubleDipper 72.9 62.3 98.6 98.8 73.0 71.5 66.0 40.2
@@
Mistral-Nemo 44.0 17.2 72.6 67.0 51.0 28.9 59.7 11.9
   Zero-shot + Evidence Retrieval 46.7 29.1 59.8 67.2 51.0 39.8 60.6 19.4
   DoubleDipper 53.3 38.7 58.0 81.0 51.4 51.9 62.9 29.5
@@
Table 2 presents the QA performance of the baseline, Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set. For brevity, we report the results of eight models here and four models in Appendix C in Table 7, which show similar trends. The results first show that prompting models to explicitly identify the relevant paragraphs before generating the answer (Zero-shot + Evidence Retrieval) leads to a performance improvement of 9.7 points on average across models over the vanilla baseline. DoubleDipper, leveraging demonstrations generated with the efficient Gemma 2B parameters, offers an additional substantial boost of 6.3 points for all models on average, culminating in an overall improvement of 16 absolute points over the vanilla baseline. Notably, while DoubleDipper produces simple QAs answerable from a single paragraph, it always surpasses the baseline in multi-hop QA datasets (HotPotQA, 2Wiki and MuSique). Likewise, DoubleDipper outperforms the baseline also on most FLenQA datasets (PIR, MonoRel and SRT), which involve synthetic True/False questions although the demonstrations in DoubleDipper are typically simple factoid questions.
While DoubleDipper exhibits strong performance overall, we observe nuanced behavior on the SRT dataset, with performance gains varying across models (improvement for Gemini, Llama 3.1 8B, Mistral 7B and Mistral Nemo). This discrepancy is likely due to the fact that SRT demands a specific type of reasoning, where models must reason over both a rule (e.g., “If X is big and X is good then X is tall”) and dispersed facts (e.g., “Erin is Good” and “Erin is furry”) to determine whether a statement (e.g., “Erin is tall”) can be derived from the context. Finally, the Lost dataset highlights a specific characteristic of DoubleDipper: while the baseline’s known positional bias (Liu et al., 2024a) masks the average improvement for some models, DoubleDipper substantially boost performance when relevant information appear in the middle of the context (further elaborated in Result 3), demonstrating its efficacy in mitigating positional biases in long-context settings.
@@
Gemma 2 9B Zero-shot + Evidence Retrieval 61.9 76.7 69.3 60.3 43.2 76.5 51.5 55.9
@@
Gemma 2 27B Zero-shot + Evidence Retrieval 85.1 96.2 97.9 96.1 84.1 90.7 53.4 77.3
DoubleDipper 83.8 97.5 85.4 97.6 74.9 93.0 59.7 78.8
@@
Llama 3.1 70B Zero-shot + Evidence Retrieval 85.1 98.2 98 89.2 82.3 93.9 52.5 81.6
DoubleDipper 87.8 98.6 100 92.8 83.8 96.5 61.6 81.6
@@
Mistral-Nemo Zero-shot + Evidence Retrieval 69.7 91.1 85.0 76.9 35.9 88.8 39.8 70.3
DoubleDipper 76.1 95.9 95.1 81.5 39.0 93.7 54.8 73.0
@@
Result 2: Learning to retrieve the evidence(s) with DoubleDipper is more effective in commercial and large open source models.
Table 3 presents the performance of the supporting paragraphs prediction for the Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set. For all commercial models, Llama 3.1 70B and the recent Mistral-Nemo-Instruct-2407, DoubleDipper predicts better the supporting paragraphs than in the zero-shot setting (+2.6 F1 for Gemini Pro, +3.4 F1 for Gemini Flash, +2.7 F1 for Llama 3.1 70B and +6.4 F1 for Mistral-Nemo-Instruct-2407). Conversely, DoubleDipper slightly hurts the performance of common open source models (e.g., -2.8 F1 for Mistral, -0.3 F1 for Llama 3.1, -1.3 F1 for Gemma 27B, etc.).
This discrepancy appears to stem from shortcut learning (Tang et al., 2023). Indeed, our demonstrations in DoubleDipper use a single evidence paragraph, and smaller models tend to overfit to this pattern, learning to retrieve only one passage even when multiple are needed. For example, under DoubleDipper, Gemma 2 9B retrieves only 1.2 paragraphs on average, compared to 2 in the “unconstrained” zero-shot setting. Larger models like Gemini Pro do not exhibit this behavior, correctly generalizing to predict an average of 2 evidence paragraphs even with single-paragraph demonstrations. This hypothesis is strongly supported by the “Lost” dataset, which requires only a single evidence paragraph. On this dataset, nearly all models, including smaller ones, benefit substantially from DoubleDipper (e.g., +12.2 F1 for Mistral 7B). Crucially, despite their suboptimal retrieval performance on multi-evidence tasks, these smaller models still produce better final answers with DoubleDipper than the zero-shot baseline, highlighting the net benefit of the approach.
@@
Following Liu et al. (2023b), Figure 3 shows the performance of Gemma 2 9B, Mixtral 8x7B and Gemini Flash for both the baseline and DoubleDipper on our sample of the Lost-of-the-middle dataset, according to the position of the document that contains the answer. See Appendix D for the performance curve of the other tested models, which show similar trends.
@@
Figure 3: Performance (accuracy) of Gemma 2 9B, Mixtral 8x7B and Gemini Flash with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023b) according to the position of the document that contains the answer.
@@ 6 Ablation Studies
Table 4: Average performance on our evaluation set with various numbers of self-generated few shot demonstrations ($k$) in DoubleDipper. See Appendix E.1 for the results on each evaluation dataset.
How many examples are needed?
In Table 4, we explore for some models the impact of varying $k$, the number of self-generated few-shot examples in DoubleDipper to 1, 3, 5, and 10. On average, a single demonstration already provides an improvement over the baseline. Three demonstrations adds another boost of 2 points, while increasing the number of demonstrations to 5 and 10 leads to a marginal improvement. This finding is in line with previous work (Brown et al., 2020a; Min et al., 2022). We conclude that a small number of examples carries most of the benefit with our method, but given additional computation budget, adding more examples does carry additional minor benefit.
DoubleDipper without identification of supporting paragraphs
To ablate the second core principle of DoubleDipper—the explicit identification of supporting paragraphs prior to answer generation—we prompt open-source models with self-generated few-shot examples consisting solely of question-answer pairs, without instructing the model to retrieve the relevant passage(s). These demonstrations may undermine DoubleDipper’s objective by encouraging models to produce answers without grounding them in the source text. The results confirms our hypothesis: removing evidence identification consistently degrades QA performance compared to the full DoubleDipper approach. When averaging results across models and datasets, this omission leads to a substantial performance drop from 54.8 to 46.6. Detailed results are provided in Appendix E.3, Table 10.
Gemma 2B Self Gemini Pro
Gemini Pro 70.4 71.6 71.6
Gemini Flash 66.1 67.5 68.1
Gemini Nano 62.1 61.7 62.8
Gemma 2B 49.5 49.5 51.0
Gemma 9B 61.2 62.0 63.5
Llama 3.1 59.9 60.1 61.5
Mistral v0.3 51.0 49.8 52.6
Mixtral 52.2 49.8 54.4
Table 5: Average performance of DoubleDipper with different models for generating the demonstrations. See Appendix E.2 for the results on each evaluation dataset.
@@
To understand the impact of the default chosen model (Gemma 2 2B) for generating the demonstrations, we conducted two additional experiments. The first experiment is Self in which we use the same model for generating the demonstrations and for answering the original question. In the second experiment, we generate the demonstrations with the best LLM used in our experiments, namely Gemini Pro. The average results are reported in Table 5 and the performance for each evaluation dataset is presented in Appendix E.2. The results show that generating the demonstrations with Gemma 2 or Self achieves similar performance, while Gemini Pro leads to a consistent increase in performance across models, indicating that future better models can improve further the performance. Please refer to Appendix E.2 for additional ablations on the few-shot generation.
Baseline ICL DoubleDipper
Gemma 2 9B 44.0 51.0 61.2
Gemma 2 27B 48.0 54.9 64.2
Llama 3.1 8B 37.2 40.7 59.9
Llama 3.1 70B 67.5 65.4 72.9
Mistral v0.3 37.4 42.0 51.0
Mixtral v0.1 42.6 45.1 52.2
GPT 4o mini 51.3 56.1 60.8
Table 6: Comparison of traditional In-Context Learning (ICL) where each demonstration example comprises a full text, a question and an answer from an external dataset to DoubleDipper where the demonstrations contain only question-answer pairs, automatically generated on the same input text.
DoubleDipper vs. Traditional ICL
Another alternative to use ICL in practice is to preprend each QA prompt by a fixed set of QA demonstrations, each composed of a context, a question and an answer. Although the demonstrations are not necessarily from the same distribution as each user query, this common practice is helpful for task recognition and for an overview of the overall format (Min et al., 2022; Pan et al., 2023). For the demonstrations, we randomly selected 3 examples from the SQuAD 2.0 dataset (Rajpurkar et al., 2018). We compare the average results of the baseline, ICL and DoubleDipper in Table 6 and report full results in Appendix E.4 in Table 11. While ICL is effective and outperforms the baseline, DoubleDipper provides an additional performance boost of 9.5 points on average.
@@ 7 Conclusion
We introduce DoubleDipper, a simple method for enhancing the performance of LLMs with long context. By recycling the input context to generate the demonstrations, DoubleDipper successfully addresses the practical challenges of ICL with long-context and outperforms multiple baselines in various QA settings, including distractor passages in the input, True/False questions and multi-hop QA.
@@ 8 Limitations
Our work has several limitations.
First, our current work focuses on multiple variants of QA tasks (distractor passages, True/False and multi-hop), where the demonstrations teach the models to identify specific evidence paragraph(s) and extract answers. Future research can extend our work to other QA settings (e.g., information seeking) and additional tasks (e.g., summarization).
Second, our evaluation set is constrained to instances that are solely in English and range between 1,000 to 4,000 tokens. While this demonstrates the method’s effectiveness, its scalability and performance on much longer contexts (e.g., 100k+ tokens) and in multilingual settings remain open questions.
Finally, while DoubleDipper is significantly more token-efficient than traditional ICL, the initial step of generating demonstrations introduces a computational and latency overhead compared to a zero-shot baseline. This presents a trade-off between inference cost and the substantial performance gains our method provides.
Appendix A Appendix
Appendix B Prompts
@@
Appendix C Additional Results
Table 7 shows the QA performance of the baseline, Zero-shot + Evidence Retrieval and DoubleDipper on our evaluation set when applied to Gemma 2B, Gemini Nano and GPT-4o-mini. Her
Avg. 2Wiki MonoRel PIR SRT HotPotQA Lost MuSique
Gemini Nano (vanilla) 41.6 10.8 72.2 66.8 55.4 21.3 59.6 5.2
   Zero-shot + Evidence Retrieval 56.5 32.0 82.4 82.4 56.4 56.7 60.5 25.2
   DoubleDipper 62.1 40.6 86.6 95.4 56.2 65.1 60.5 30.4
GPT-4o-mini (vanilla) 51.3 19.0 78.6 89.6 71.6 23.3 67.8 9.3
   Zero-shot + Evidence Retrieval 53.8 19.4 89.6 93.0 63.2 26.2 67.9 17.6
   DoubleDipper 60.8 29.8 94.8 96.4 61.6 53.7 64.7 24.5
Gemma 2 2B (v2) (vanilla) 38.6 8.9 71.8 68.6 51.2 13.3 49.6 6.5
   Zero-shot + Evidence Retrieval 42.0 22.3 66.8 70.6 40.2 30.6 47.6 16.2
   DoubleDipper 49.5 23.7 85.8 81.6 50.0 39.9 46.7 18.8
Mixtral 7x8B (v0.1) (vanilla) 42.6 13.7 73.0 66.2 51.0 18.2 67.7 8.4
   Zero-shot + Evidence Retrieval 47.4 18.8 81.8 73.6 50.6 26.3 67.9 13.1
   DoubleDipper 52.2 22.3 91.8 86.0 47.8 35.1 66.6 16.0
Appendix D Lost-in-the-middle
@@
Figure 5: Performance (accuracy) of Gemma 2 2B, Mistral 7B, Gemini Nano and Gemini Pro with and without DoubleDipper on our sample of the Lost-in-the-middle dataset (Liu et al., 2023b) according to the position of the document that contains the answer.
Appendix E Analysis
E.1 Impact of the Number of Demonstrations in DoubleDipper
@@
E.2 Impact of the few-shot generator
@@
Can we generate demonstrations with smaller LLMs?
We next study whether smaller, more efficient models could be used to generate demonstrations for DoubleDipper. To do this, we created demonstrations using two smaller LLMs, Gemma 3 1B (Kamath et al., 2025) and Qwen 2.5 0.5B (Bai et al., 2023a), and provid them to a subset of our evaluation models (Llama 3.1, Gemma 2 9B, Gemma 2 27B, and Mistral Nemo). The results show that DoubleDipper maintains its advantage over baselines, exhibiting only a modest performance drop of less than 2 points compared to using demonstrations from the larger Gemma 2 2B. This finding suggests that DoubleDipper remains effective even when its demonstration generation component is replaced with more lightweight models.
Can DoubleDipper benefit from incorrect demonstrations?
To answer this question, we prompt Gemma 2 2B to generate a question and an incorrect answer and provide these demonstrations to a sample of our tested models (Llama 3.1, Gemma 2 9B, Gemma 2 27B and Mistral Nemo). The results are mixed and not conclusive: some models barely benefit or suffer from incorrect demonstrations (+1.6 for Gemma 2 9B, -0.4 for Gemma 2 27B), while others somehow benefit from incorrect demonstrations (+4.8 for Llama 3.1 8B and +5.1 for Mistral Nemo over the baseline). Critically, however, all models still perform substantially worse than DoubleDipper with correct demonstrations. A possible explanation for the unexpected gains is that even incorrect examples provide useful structural guidance for the task format, a phenomenon observed in prior work (Min et al., 2022).
E.3 Impact of the identification of supporting paragraphs in the QA generation